OpenStudy (dylanfwlr):

Find an equation of a parabola with a vertex at the origin and directrix y=-2.5

OpenStudy (dylanfwlr):

@amistre64

OpenStudy (amistre64):

i always have to do this the long way ....

OpenStudy (dylanfwlr):

i dont know any way to do it :/

OpenStudy (amistre64):

y = ax^2 since the vertex is at the origin ... we would need to define 1 point to solve for a

OpenStudy (amistre64):

one thing I do know is that a point that is directly across from the focus has a y value equal to the distance from vertex to directrix; and an x value that is twice that distance

OpenStudy (amistre64):

so we need to solve for a knowing that the point (5, 2.5) has to be on the parabola

OpenStudy (dylanfwlr):

alright

OpenStudy (amistre64):

a = y/x^2 a = 2.5/25 what is the value for a?

OpenStudy (dylanfwlr):

1

OpenStudy (amistre64):

i was thinking .1

OpenStudy (dylanfwlr):

oh sorry i though 25 was 2.5

OpenStudy (amistre64):

yeah, its hard for these old eyes of mine to see those dots too :)

OpenStudy (amistre64):

so, y = .1 x^2 should do it

OpenStudy (dylanfwlr):

that's it?

OpenStudy (amistre64):

http://www.wolframalpha.com/input/?i=y%3D.1x%5E2 the wolf likes it

OpenStudy (dylanfwlr):

cool. (i had to change it to a fraction so i just put 1/10)

OpenStudy (amistre64):

thats fine too

OpenStudy (amistre64):

the key is, know where the vertex is, and find that point that is directly across from the focal point. thats how i do it

OpenStudy (dylanfwlr):

there is a few more but i'm going to try them myself before i post

OpenStudy (amistre64):

good luck ;)